OpenStudy (rainbow_dash):

After a gas-filled balloon is released, it rises 90 feet in the first minute. In the second minute, the balloon rises 120 feet, and in the third minute, it rises 150 feet. How many feet would the balloon rise in 8 minutes?

OpenStudy (anonymous):

the answer is 300

OpenStudy (rainbow_dash):

thank you

OpenStudy (anonymous):

@jasonlam Just don't give the answer next time.

OpenStudy (australopithecus):

90, 120, 150 so we have points where x = minutes, y = feet (1,90) (2,120) (3,150) Use the equation of the line y= mx + b where m = slope b = y-intercept use the equation m = (y1-y2)/(x1-x2) I'm using points (2,120) and (3,150) 120 - 90/3-2 = 30/1 = 30 = m y = mx + b plug in m y = (30)x + b solve for b by plugging in a point I used (1, 90) y = 30x + b 90 = 30(1) + b 90-30 = b 60 = b b = 30 so we have the equation y = 30x + 60 checking the answer y = 30(1) + 60 y = 90 y = 30(2) + 60 y = 120 y = 30(3) + 60 y = 150 plug 8 into the equation, to find the amount of feet

OpenStudy (australopithecus):

hope this helps although I think you are done with this question ha